OpenStudy (anonymous):

What is the value of sin22.5

OpenStudy (anonymous):

you would probably need a calculator for finding cosine, sine, and tangent, but the answer is 0.38

OpenStudy (unklerhaukus):

The Half Angle Formula for sine is:\[\boxed{~\\~ \sin^2\theta=\dfrac{1-\cos(2\theta)}2~\\~}\]

OpenStudy (unklerhaukus):

so that \[\sin\theta=\sqrt{\dfrac{1-\cos(2\theta)}2}\]

OpenStudy (unklerhaukus):

note that 22.5*2 = 45

OpenStudy (unklerhaukus):

@sangineni come back . . .